# Dispersive regime of multiphoton qubit-oscillator interactions

### v1

@@ III.2 Non-RWA dispersive regime
$\hat{H}_{\text{Disp}}^{(n)}$ $=\hat{U}_{\text{Disp}}^{(n)\dagger}\hat{H}_{n-\text{R}}\hat{U}_{\text{Disp}}^{(n)}$
$=\hat{H}_{n-\text{R}}+[\hat{H}_{n-\text{R}},\lambda_{n}\hat{X}_{-n}+\overline{\lambda}_{n}\hat{Y}_{-n}]$
$\,\,\,\,\,\,\,+\frac{1}{2}\Big{[}[\hat{H}_{n-\text{R}},\lambda_{n}\hat{X}_{-n}+\overline{\lambda}_{n}\hat{Y}_{-n}],$
@@
where $\xi_{n}=g_{n}^{2}/\Sigma_{n}$ is the $n$-photon Bloch-Siegert shift coefficient. The non-RWA regime transformation yields an additional shift, $\xi_{n}$, dependent on the counter-rotating terms which in turn oscillate with $\omega_{q}+n\omega_{o}$. The commutators $[\hat{X}_{+n},\hat{Y}_{-n}],\,[\hat{Y}_{+n},\hat{X}_{-n}],$ and $[\hat{Y}_{-n},\hat{Y}_{+n}]$ give rise to the same combinatorial coefficients as $[\hat{X}_{+n},\hat{X}_{-n}]$ and are also discussed in App. A. The non-RWA regime Hamiltonian exhibits a generalized $2n$-photon squeezing term, $\hat{\sigma}_{z}(\hat{a}^{\dagger 2n}+\hat{a}^{2n})$. This is consistent with the one-photon dispersive calculations presented in Sec II. Unlike in the RWA regime, here, the qubit-state-independent oscillator anharmonicity is negative and pushes the energy levels down.
@@
Figure 2: Multiphoton qubit-oscillator interaction spectra in the dispersive regime. The analytic spectra from the RWA and non-RWA perturbative expansions are compared with the numerically-obtained spectra of Eq. (III) with $\eta=0.01$. For each plot, $\Delta_{n}$ is fixed, and, as such, the $x$-axis ($g_{n}$) can be thought of as tuning the perturbation parametrs $\lambda_{n}$ and $\overline{\lambda}_{n}$. The analytic results, in general, deviate at smaller couplings strengths, or, equivalently, perturbation parameters, for higher-order interactions (see (a/b) compared to (c/d)). The non-RWA spectra deviates at smaller coupling strengths than its RWA counter-part. Additionally, the higher-energy levels RWA and non-RWA analytic spectra deviate at smaller coupling strengths.
@@ III.3 Comparison with numerical results
We now compare the analytically derived results of the RWA and non-RWA spectra with the numerically-obtained spectra of the $n$R model [Eq. (III) with $\eta=0.01$]. We prelude the discussion with a reminder of the looming spectral instabilities for higher orders of $n$. The spectral instabilities of the nR model for $n>2$ restrict the low-energy regime in a smaller set of parameters than $n=2$. Thus, the low-energy stable regime shrinks for higher-order interactions. This means that, in general, the regime of validity for the perturbation theory becomes smaller for larger $n$. Additionally, the dispersive regime where the qubit-oscillator system can be approximately described in its bare basis comprises a smaller range of parameters for higher orders of $n$. Another issue to keep in mind is the fact that the usual non-RWA effects become non-negligible for larger $g_{n}/\omega_{o}$ — at least in the usual linear case. This suggests the need for a spectral classification of the $n$R model that is distinct from the spectral classification of the typical linear Rabi model as in Refs. [57, 58].
@@
In Fig. 2, we compare the analytical results from second-order SW perturbation theory with numerical diagonalization of the $n$R Hamiltonian [^2] In the case of $n=2$, the numerically-obtained spectra match the quasi-analytic solutions of Refs. [16, 17]. We start by noting general trends. The description provided by perturbation theory significantly improves when $\lambda_{n}$ (and $\overline{\lambda}_{n}$) gets smaller, as expected. This is seen in the improvement of overlap between the numerical and analytical spectra between Fig. 2(a) and (b) for $n=2$, and (c) and (d) for $n=3$, where the ranges of $\lambda_{n}$ (and $\overline{\lambda}_{n}$) become smaller for higher $|\Delta_{n}|.$ The deviation from the analytic description occurs in a similar fashion to the deviation of the general $n$R spectra from the $n$JC spectra in two ways. The first trend is one within a fixed interaction order; the higher energy levels deviate at smaller ratios of $g_{n}/\omega_{o}$. The second trend is that the higher-order interactions deviate at smaller ratios of $g_{n}/\omega$ — this can be immediately attributed to the smaller dispersive regime for higher-order interactions as discussed earlier. Figures 2(c) and (d) show the dispersive regime for $n=3$, and we observe the two trends noted above; locally for a fixed interaction order, the higher energy levels deviate at smaller coupling strengths, and globally when comparing (a) and (b) to (c) and (d), the higher-order interactions comprise a smaller range of parameters in the dispersive regime.
A surprising occurrence is the deviation of the non-RWA analytical spectra from the numerical calculations in many instances at smaller coupling strengths than the RWA spectra. In other words, the non-RWA results are a worse approximation than the RWA results. This can be seen in Fig. 2. This suggests that the role of the counter-rotating terms in $n$R model is not well-represented in the non-RWA SW transformation. Thus, a simple extension of the generators in the generalized multiphoton SW transformation is not enough to extend the validity of the perturbative expansion spectra. This is very different from the linear interaction where the non-RWA generators extend the validity of the perturbative expansion and analytic spectra to regimes of larger coupling strengths [29, 30].
Figure 3: Effective qubit frequency due differing interaction orders as a function of oscillator coherent state amplitude The ratio of effective qubit frequency to bare qubit frequency is plotted as a function of the amplitude of the coherent state populating the oscillator. The black, purple and pink lines correspond to $n=1$, $n=2$ and $n=3$ dispersive interactions, respectively; solid lines represent $g_{n}=0.02\omega_{o}$ and dashed lines represent $g_{n}=0.03\omega_{o}$.
@@ III.4 Experimental signature of dispersive multiphoton qubit-oscillator interactions
As the RWA perturbation theory proved to be more effective than its non-RWA counterpart, we rely on the Hamiltonian of Eq. (III.1). We consider the effective (dressed) qubit frequency as a function of the oscillator population and its powers,
@@ IV Multiple qubits coupled to a single oscillator
The spectral collapse for $n=2$ persists in the multiqubit case. The presence of multiple qubits changes the critical coupling at which the collapse occurs. It was shown for the case of $n=2$ that the collapse occurs at a smaller collective critical collapse coupling [24]. Explicitly, for a fixed oscillator frequency, if we take a single qubit-oscillator system where the critical collapse coupling is $g_{2,c}$, the individual qubit couplings at which the multipartite system’s spectral collapse occurs are all less than the single qubit case, $g_{2}^{(1)},...,g_{2}^{(N)}<g_{2,c}$; the shifted multipartite ($N$ qubit) critical collapse coupling $\widetilde{g}_{2,c}^{N}$ is less than $g_{2,c}$ due to coupling enhancement in Dicke-type models. When all individual couplings are the same, the enhancement scales as $\sqrt{N}$. For varying individual coupling strengths, there is no simple scaling factor describing the enhancement of the individual coupling strengths. However, some approximate relations can be derived, for example using a mean-field treatment of the interactions between the different subsystems [60]. Additionally, the discussion on the spectral instabilities for $n>2$ follows from the single qubit-oscillator case and applies just as well. Figure 4 shows the $n$D Hamiltonian spectra (in red) for a two-qubit system ($N=2$). Here, we detune one qubit above the oscillator frequency and the other below it where both possess an $n$-photon detuning of $|\Delta_{n}^{(1)}|=|\Delta_{n}^{(2)}|=\omega_{o}.$ For a single plot in Fig. 4, the qubits and oscillator frequencies are fixed as well as one of the qubits’ coupling strengths. Then, we can see the effect of changing the other qubit’s coupling strength on the spectra in general and the spectral collapse in particular.
@@ IV.2 Non-RWA dispersive regime
Figure 4: Spectra of the two-qubit two-photon Dicke model contrasted with petrubative spectra. We assume $|\Delta_{2}^{(j)}|=\omega_{o}$; one qubit is detuned above the oscillator frequency and the other is detuned below. In this case, $\widetilde{\chi}_{2}^{(1,2)}=0.$ Solid lines represent numerical results, dashed lines represent RWA analytical results, and dotted lines represent non-RWA analytical results. For each plot we fix the two qubit frequencies, oscillator frequency and one of the qubits’ coupling strengths. The coupling strength of the other qubit is varied, and the spectra are plotted as a function of it. The spectral collapse shifts to smaller critical coupling ratios which occurs due to the collective qubit cooperative effects. The RWA results generally provide an accurate description in the low collective coupling regimes, but the non-RWA predictions deviate more rapidly from the numerical spectra. For higher collective coupling (sum of individual couplings), the accuracy of perturbative results becomes less accurate (see main text for detailed commentary).
@@ IV.3 Comparison with numerical results
It is possible to exactly diagonalize the RWA Hamiltonian of Eq. (IV.1) by simply diagonalizing the two-qubit system. Essentially, the Hamiltonian becomes a block-diagonal matrix of $4\times 4$ blocks. Similarly, in the non-RWA case, we can first apply the extended transformation accounting for the $2n$-photon squeezing term presented in Eq. (33), and then diagonalize the two-qubit part of the Hamiltonian. We now compare the numerically obtained spectra to those analytically obtained from the RWA and non-RWA Hamiltonians.
The analytical spectra produced by the RWA and non-RWA expansions are contrasted with the numerical spectra of the two-qubit two-photon Dicke Hamiltonian in Fig. 4. For a single plot, we fix one qubit’s coupling to the oscillator and we vary the other qubit’s coupling. Here, we assume $\Delta_{2}^{(1)}=-\Delta_{2}^{(2)}=\omega_{o}.$ It is important to note that in this case, the RWA Hamiltonian predicts the vanishing of the mediated photon-number-enhanced qubit-qubit interaction as $\widetilde{\chi}_{2}^{(1,2)}=0$. Meanwhile, the non-RWA Hamiltonian predicts its persistence as $\widetilde{\xi}_{2}^{(1,2)}\neq 0$, which should significantly affect the resulting spectra. Figure 4(a) shows the regime where the fixed qubit coupling is well in its single qubit-oscillator dispersive regime, i.e. $g_{2}^{(2)}\ll g_{2,c}$ with $g_{2,c}$ being the single qubit-oscillator two-photon critical collapse coupling. In this regime, the collective qubit cooperative effects is relatively weak which only slightly shifts the aforementioned collective critical collapse coupling, $\widetilde{g}_{n,c}^{N}$ ($N=n=2$), slightly lower than $g_{n,c}.$ We find that the perturbative spectra shows very similar trends to the single qubit-oscillator case; the lower-energy states better match the numerical results, and the RWA expansion generally better approximates the system whereas the non-RWA predictions deviate at lower coupling strengths. When both qubits’ couplings are in the single qubit-oscillator dispersive regime, i.e. $g_{2}^{(j)}\ll|\Delta_{2}|^{(j)},\,|\Delta_{2}|^{(2)},\,g_{2,c}$ ($j=1,2$), the perturbative description of the system is in excellent match with the numerics. However, when the one of the qubits is outside its individual regime of validity, the collective critical collapse coupling shifts more significantly as shown in Figs. 4(b) and (c). This also results in the deviation of the perturbative spectra and, so, they become valid in for a smaller range of parameters. The non-RWA spectra significantly deviate in the case higher-energy states in Fig. 4(b). This is because the non-RWA Hamiltonian’s qubit-qubit interaction term does not vanish and with growing qubit coupling strengths and higher Fock states–hence higher energy states– becomes a significant contribution. This reaffirms the conclusion of the single qubit-oscillator results where the non-RWA extension of the SW transformation becomes a less accurate description particularly, for higher-energy states and generally, at lower coupling strengths.
@@ IV.4 Oscillator-mediated tunably coupled multiqubit system
The photon-number dependence of the mediated qubit-qubit interaction allows for the tuning of the effective qubit frequencies and mediated qubit-qubit coupling by carefully selecting the oscillator state. As an example, we can take two qubits interacting with an oscillator via an $n$-photon interaction. Since the RWA expansion results in a better approximation than the non-RWA case, we consider the RWA Hamiltonian of Eq. (IV.1). Similarly to the calculations carried out in Sec. III.4, we consider the oscillator to be populated by a coherent state |$\ket{\alpha}.$ We now consider the effective two-qubit Hamiltonian. Then, the effective Hamiltonian (ignoring constant offsets) is

### v2

@@ III.2 Non-RWA dispersive regime
$\hat{H}_{\text{Disp}}^{(n)}$ $=\hat{U}_{\text{Disp}}^{(n)\dagger}\hat{H}_{n\text{R}}\hat{U}_{\text{Disp}}^{(n)}$
$=\hat{H}_{n\text{R}}+[\hat{H}_{n\text{R}},\lambda_{n}\hat{X}_{-n}+\overline{\lambda}_{n}\hat{Y}_{-n}]$
$\,\,\,\,\,\,\,+\frac{1}{2}\Big{[}[\hat{H}_{n\text{R}},\lambda_{n}\hat{X}_{-n}+\overline{\lambda}_{n}\hat{Y}_{-n}],$
@@
where $\xi_{n}=g_{n}^{2}/\Sigma_{n}$ is the $n$-photon Bloch-Siegert shift coefficient. The non-RWA regime transformation yields an additional shift, $\xi_{n}$, dependent on the counter-rotating terms which oscillate with $\omega_{q}+n\omega_{o}$. The commutators $[\hat{X}_{+n},\hat{Y}_{-n}],\,[\hat{Y}_{+n},\hat{X}_{-n}],$ and $[\hat{Y}_{-n},\hat{Y}_{+n}]$ give rise to the same combinatorial coefficients as $[\hat{X}_{+n},\hat{X}_{-n}]$ and are also discussed in App. A. The non-RWA regime Hamiltonian exhibits a generalized $2n$-photon squeezing term, $\hat{\sigma}_{z}(\hat{a}^{\dagger 2n}+\hat{a}^{2n})$. This is consistent with the one-photon dispersive calculations presented in Sec II. Unlike in the RWA regime, here, the qubit-state-independent oscillator anharmonicity is negative and pushes the energy levels down.
@@
Figure 3: Multiphoton qubit-oscillator interaction spectra in the dispersive regime. The analytic spectra from the RWA and non-RWA perturbative expansions are compared with the numerically-obtained spectra. For each plot, $\Delta_{n}$ is fixed (which also fixes $\Sigma_{n}$), and, as such, the x-axis ($g_{n}$) can be thought of as tuning the perturbation parameters $\lambda_{n}$ and $\overline{\lambda}_{n}$. The top row ((a)-(d)) corresponds to $n=2$, and the bottom row ((e)-(h)) corresponds to $n=3$. Plots (a), (b), (e) and (f) are in the RWA regime where $\Delta_{n}\ll\Sigma_{n}$. Similarly, plots (c), (d), (g) and (h) are in the non-RWA regime where $\Delta_{n}\sim\Sigma_{n}.$ For smaller values of $\Delta_{n}$, the perturbative spectra deviate from the numerical ones at smaller coupling strengths. This is because smaller $\Delta_{n}$ values correspond to larger $\lambda_{n}$ values. When $\Delta_{n}\sim\Sigma_{n}$, the non-RWA spectra become a better approximation of the numerical results than the RWA spectra.
@@ III.3 Comparison with numerical results
We now compare the analytically derived results of the RWA and non-RWA spectra with the numerically-obtained spectra of the $n$R model (for details on the numerical spectra and stabilization, see App. B).
We prelude the discussion with a reminder of the looming spectral instabilities for higher-order interactions. The spectral instabilities of the nR model for $n>2$ restrict the low-energy regime in a smaller set of parameters than $n=2$. Thus, the low-energy stable regime shrinks for higher-order interactions. This means that, in general, the regime of validity of perturbation theory becomes smaller for larger $n$. Additionally, the dispersive regime where the qubit-oscillator system can be approximately described in its bare basis comprises a smaller range of parameters for higher order $n$. In addition to the regime of validity set by the size of the perturbation parameters, one can derive a heuristic bound based on a critical photon number in a similar fashion to the one-photon critical photon number discussed in Sec.II. The effect of SW perturbation theory in the RWA regime can be described as a linearization of the square root term in the $n$JC energy eigenvalues of Eq. (24). Let $n_{\text{ph}}$ be the oscillator occupation. When $g_{n}\sqrt{(n_{\text{ph}}+n)!/n_{\text{ph}}!}\sim\Delta_{n}$, the approximate linearization of the square root begins to break down. Thus, using the fact that $\sqrt{(n_{\text{ph}}+n)!/n_{\text{ph}}!}$ scales as $n_{\text{ph}}^{n/2}$, we can arrive at a heuristic critical photon number for the $n$JC dispersive regime; $n_{\text{ph},c}^{(n)}\propto(\Delta_{n}/g_{n})^{2/n}.$ This heuristic bound predicts a shrinking regime of validity for the RWA dispersive spectra for higher order $n.$ Figure 2 shows the behaviour of the critical photon number in the cases of $n=1,2,3$ and $4$ as a function of $g_{n}$ for different values of $\Delta_{n}.$
In Fig. 3, we compare the analytical results from second-order SW perturbation theory with numerical diagonalization of the $n$R Hamiltonian [^2] In the case of $n=2$, the numerically-obtained spectra match the quasi-analytic solutions of Refs. [16, 17].. Here, we use an oscillator truncation of $N_{T}=300$ and focus on the states that correspond to the first few bare qubit-oscillator states at $g_{n}=0$, namely, the first few states |$\ket{g/e}\ket{l}$. The details regarding stabilization and justification for using $N_{T}=300$ are found in App. B.
@@
We start by noting general trends. The description provided by perturbation theory is most accurate when $\lambda_{n}$ is small for the RWA regime (and, additionally, when $\overline{\lambda}_{n}$ is small for the non-RWA regime), as expected. This is seen in the improvement of overlap between the numerical and RWA analytical spectra between Fig. 3(a) and (b) for $n=2$, and (e) and (f) for $n=3$. For each plot $\Delta_{n}$ is fixed which makes the x-axis, $g_{n}/\omega_{o}$ proportional to $\lambda_{n}$. Thus, for plots with larger $\Delta_{n}$, the range of $\lambda_{n}$ values used in the plot is smaller. The deviation from the analytic description occurs in a similar fashion to the deviation of the general $n$R spectra from the $n$JC spectra in two ways. The first trend is one within a fixed interaction order; the higher energy levels deviate at smaller ratios of $g_{n}/\omega_{o}$. The second trend is that the higher-order interactions deviate at smaller ratios of $g_{n}/\omega$ — this can be immediately attributed to the smaller dispersive regime for higher-order interactions as discussed earlier. Figures 3(c) and (d) show the non-RWA dispersive regime for $n=2$. We observe that as $|\Delta_{n}|$ becomes of comparable magnitude to $\Sigma_{n}$, the non-RWA analytic spectra (black dotted lines) serve as better approximations for the numerical spectra than their RWA counterparts. This is also the case for $n=3$ as seen in Fig. 3(g) and (h).
[…]
@@ III.4 Experimental signature of dispersive multiphoton qubit-oscillator interactions
Situating ourselves in the RWA regime (what follows works just as well in the non-RWA regime), we rely on the Hamiltonian of Eq. (III.1). We consider the effective (dressed) qubit frequency as a function of the oscillator population and its powers,
@@
Figure 4: Fidelity of time-evolved states under the dispersive approximation. The qubit and oscillator subsystem fidelities are obtained by time-evolving the initial states under Eq. (III.1) and Eq. (20) for the case of $n=2$ then tracing out the oscillator and qubit, respectively.
@@ IV Multiple qubits coupled to a single oscillator
Figure 5: Effective qubit frequency due differing interaction orders as a function of oscillator coherent state amplitude. The ratio of effective qubit frequency to bare qubit frequency is plotted as a function of the amplitude of the coherent state populating the oscillator. The black, purple and pink lines correspond to $n=1$, $n=2$ and $n=3$ dispersive interactions, respectively; solid lines represent $g_{n}=0.02\omega_{o}$ and dashed lines represent $g_{n}=0.03\omega_{o}$.
@@
The spectral collapse for $n=2$ persists in the multiqubit case. The presence of multiple qubits changes the critical coupling at which the collapse occurs. It was shown for the case of $n=2$ that the collapse occurs at a smaller collective critical collapse coupling [24]. Explicitly, for a fixed oscillator frequency, if we take a single qubit-oscillator system where the critical collapse coupling is $g_{2,c}$, the individual qubit couplings at which the multipartite system’s spectral collapse occurs are all less than the single qubit case, $g_{2}^{(1)},...,g_{2}^{(N)}<g_{2,c}$; the shifted multipartite ($N$ qubit) critical collapse coupling $\widetilde{g}_{2,c}^{N}$ is less than $g_{2,c}$ due to coupling enhancement in Dicke-type models. When all individual couplings are the same, the enhancement scales as $\sqrt{N}$. For varying individual coupling strengths, there is no simple scaling factor that describes the enhancement of individual coupling strengths. However, some approximate relations can be derived, for example using a mean-field treatment of the interactions between the different subsystems [63]. Additionally, the discussion on the spectral instabilities for $n>2$ follows from the single qubit-oscillator case and applies just as well. Similarly to the one-photon Dicke model, the multiphoton Dicke model exhibits normal and superradiant phases. The dispersive regime is concerned with qubit-oscillator coupling strengths that are well within the normal phase [64, 65].
@@ IV.3 Comparison with numerical results
[…]
$\hat{\widetilde{H}}^{(n)}_{j}=\frac{1}{2}\begin{pmatrix}\Sigma^{(1,2)}+\Upsilon^{(1,2)}_{+j,n}+\mu_{j,n}^{(1,2)}&0&0&\Xi^{(1,2)}_{j,n}\\ 0&\Delta^{(1,2)}+\Upsilon^{(1,2)}_{-j,n}+\mu_{j,n}^{(1,2)}&\Xi^{(1,2)}_{j,n}&0\\ 0&\Xi^{(1,2)}_{j,n}&-\Delta^{(1,2)}-\Upsilon^{(1,2)}_{-j,n}+\mu_{j,n}^{(1,2)}&0\\ \Xi^{(1,2)}_{j,n}&0&0&-\Sigma^{(1,2)}-\Upsilon^{(1,2)}_{+j,n}+\mu_{j,n}^{(1,2)}\end{pmatrix},$ (45)
where
$\Sigma^{(1,2)}=\omega_{q}^{(1)}+\omega_{q}^{(2)},$
$\Delta^{(1,2)}=\omega_{q}^{(1)}-\omega_{q}^{(2)},$
$\Xi^{(1,2)}_{j,n}=(\widetilde{\chi}_{n}^{(1,2)}-\widetilde{\xi}_{n}^{(1,2)})\sum_{k=1}^{n-1}C_{n,k}^{-}j^{k},$
$\Upsilon^{(1,2)}_{\pm j,n}=[(\chi^{(1)}_{n}+\xi^{(1)}_{n})\pm(\chi^{(2)}_{n}+\xi^{(2)}_{n})]\sum_{k=0}^{n}C_{n,k}^{+}j^{k},$
and
$\mu_{j,n}=2j\omega_{o}+(\chi_{n}^{(1)}+\chi_{n}^{(2)}-\xi_{n}^{(1)}-\xi_{n}^{(2)})\sum_{k=1}^{n-1}C_{n,k}^{-}j^{k}.$
The RWA form can be recovered by setting $\xi^{(j)}=0$ as well as zeroing the matrix entries |$\outerproduct{ee}{gg}$ | and |$\outerproduct{gg}{ee}$ |; this is equivalent to transforming to the appropriate rotating frame and neglecting the fast-oscillating terms.
The analytical spectra produced by the RWA and non-RWA expansions are contrasted with the numerical spectra of the two-qubit two-photon Dicke Hamiltonian in Fig. 6. Here, we use an oscillator truncation $N_{T}=300$ to produce the numerical spectra. For a single plot, we fix one qubit’s coupling to the oscillator and we vary the other qubit’s coupling. Here, we assume $\Delta_{2}^{(1)}=\Delta_{2}^{(2)}=\Delta_{2}.$ Figure 6(a) and (c) show the multiqubit RWA and non-RWA regimes, respectively, where the collective qubit cooperative effects is relatively weak which only slightly shifts the aforementioned collective critical collapse coupling, $\widetilde{g}_{n,c}^{N}$ ($N=n=2$), slightly lower than $g_{n,c}.$ We find that the perturbative spectra shows very similar trends to the single qubit-oscillator case; the lower-energy states better match the numerical results, the RWA analytical spectra are accurate while $\Delta_{2}\ll\Sigma_{2}$, and the non-RWA analytical spectra are accurate for $\Delta_{2}\sim\Sigma_{2}.$ Additionally, the impact of stronger cooperative effects on the RWA and non-RWA regimes is shown in Fig. 6(b) and (d), respectively. In this case the second qubit’s coupling to the oscillator is larger and this causes the analytical spectra to deviate at a smaller coupling of the first qubit.
@@ IV.4 Oscillator-mediated tunably coupled multiqubit system
The photon-number dependence of the mediated qubit-qubit interaction allows for the tuning of the effective qubit frequencies and mediated qubit-qubit coupling by carefully selecting the oscillator state. As an example, we can take two qubits interacting with an oscillator via an $n$-photon interaction. For simplicity, we consider the RWA Hamiltonian of Eq. (IV.1) and the discussion follows exactly for the non-RWA case. Similarly to the calculations carried out in Sec. III.4, we consider the oscillator to be populated by a coherent state |$\ket{\alpha}.$ We now consider the effective two-qubit Hamiltonian. Then, the effective Hamiltonian (ignoring constant offsets) is
@@
Figure 6: Two-qubit two-photon dispersive spectra. For each plot, the first qubit’s coupling to the oscillator, $g_{2}^{(1)}$ is varied along the x-axis while the second qubit’s coupling, $g_{2}^{(2)}$, is fixed. We set $\Delta_{2}^{(1)}=\Delta_{2}^{(2)}=0.5\omega_{o}$ for (a),(b) and $\Delta_{2}^{(1)}=\Delta_{2}^{(2)}=6\omega_{o}$ for (c),(d). The RWA and non-RWA analytics are an excellent approximation of the numerics when the cooperative effects are weak as in (a) and (c). When the cooperative effects are stronger, the analytic spectra deviate at smaller coupling values of the first qubit in (b) and (d).